### v1

@@ Abstract
Ghost imaging (GI) has demonstrated diverse imaging capabilities enabled by its encoding-decoding-based computational imaging mechanism. Accordingly, information-theoretic studies have emerged as a promising avenue for probing the fundamental performance bounds of of GI and related computational imaging paradigms. However, the design of information-theoretically optimal encoding strategies remains largely unexplored, primarily due to the intractability of the prior probability density function (PDF) of an unknown scene. Here, by leveraging the ability of recursively estimating the PDF of the object to be imaged via Bayesian filtering, we propose to establish an adaptive information-maximization encoding (AIME) design framework. Based on the adaptively estimated posterior PDF from previously acquired measurements, the expected information gain of subsequent detections is evaluated and maximized to design the corresponding encoding patterns in a closed-loop manner. Within this framework, the theoretical form of the information-optimal encoding under representative physical constraints is analytically derived. Corresponding experimental results show that, GI systems employing information-optimal encoding achieve markedly improved imaging performance compared with conventional fixed point-to-point imaging without relying on additional heuristic regularization schemes, particularly in low signal-to-noise ratio regimes. Moreover, the proposed strategy consistently enables significantly enhanced information acquisition capability compared with existing encoding strategies, leading to substantially improved imaging quality. These results establish a principled information-theoretic foundation for optimal encoding design in computational imaging paradigms, provided that the forward model can be accurately characterized.
@@ 1 Introduction
Ghost imaging (GI) [moreau2018ghost] has demonstrated a broad range of capabilities, including acquisition of high-dimensional (H-D) information through low-dimensional detection [eshun20253d], enhancement of two-dimensional resolution via H-D light-field information [tong2020breaking, tong2025single], non-locally decoupling detection efficiency and imaging resolution [yuan2021single], and reduction of the unnecessary detection redundancy [hu2019optimization]. These advantages have enabled its diverse applications in imaging LiDAR [deng2017performance, pan2021micro], single-shot H-D imaging cameras [liu2020spectral, chu2021spectral], super-resolution microscopy [li2019single, chen2025multicolor], low-dose and multi-dimensional radiographic imaging [zhang2018tabletop, he2021single], imaging in special band [stantchev2020real, bogdanov2025ghost], and task-oriented imaging [wu2025image, abbas2025target]. From the perspective of imaging information acquisition [barrett2013foundations, shensheng2022review], GI can be interpreted as an encoding-decoding-based computational imaging paradigm [hu2022ghost], where the imaging object is encoded by randomly fluctuating light fields, and the desired image is decoded through high-order light-field correlation. This formulation naturally places GI within an information-theoretic framework, under which analyses of information acquisition efficiency and fundamental performance are in principle attainable [enrong2013mutual, li2017negative, hu2021correspondence].
@@
In this Article, we explicitly investigate the possibility to adaptively establish the information-optimal encoding for GI, and propose a general adaptive information-maximization encoding (AIME) framework. It enables quantitative evaluation of the information content acquired in the encoding-detection process, and facilitates the design of optimal encoding under practical constraints through information maximization. Theoretical analysis and numerical simulations consistently show that, under a total-energy constraint corresponding to a fixed encoding illumination energy, the information-maximization encoding converges to a point-like illumination pattern. In contrast to traditional imaging, the resulting scanning order is adaptively determined by the object rather than following a deterministic sequence. Owing to this adaptive design, the proposed AIME strategy achieves superior imaging performance compared to conventional “fixed point-to-point” imaging mode without relying on additional heuristic regularization schemes, particularly in low signal-to-noise ratio (SNR) regimes. Comparison with typical encoding patterns also show that the information-optimal encoding consistently yields improved imaging quality across a wide range of sampling ratios (SRs) and SNR conditions. The proposed framework establishes a principled information-theoretic foundation for characterizing and approaching the fundamental performance of computational imaging paradigms.
@@ 2.1 Principle of AIME for GI
We formulate AIME as a closed-loop information-theoretic framework for GI, in which the encoding patterns are treated as variables to be optimized according to the evolving statistical state of the estimated imaging object. As illustrated in Fig. 1, at each measurement step, the statistical state is represented by a posterior PDF inferred from previously detected signals, based on which the expected information gain of a candidate encoding is evaluated. The subsequent encoding is then determined by maximizing the information criterion under explicit physical followed by a recursive update of the posterior PDF after measurement. Within this formulation, encoding, detection, and statistical inference in GI are integrated into a unified framework, enabling information-optimal encoding to be defined and implemented sequentially without requiring special assumption on prior knowledge of the object statistics.
Figure 1: Overview of the proposed framework for adaptive information-maximization encoding (AIME) in GI. The optimal encoding is optimized by maximizing the information of the object to be imaged contained in the detection signal. The information is calculated by the forward detection model and posterior PDFs.
@@ 2.1.1 Bayesian formulation for recursive posterior PDF tracking
To enable information-theoretic evaluation and optimization of encoding, the statistical state of the imaging object is characterized by a posterior PDF that is updated recursively as measurements are acquired. To this end, the forward encoding-detection process of GI is formulated to construct the measurement-likelihood function $p({\bf z}\mid{\bf x})$, specifically as [hu2022ghost]
@@
Then, an initial Gaussian distribution $p({\bf x}\mid{\bf Z}_{0})=p_{0}({\bf x})={\mathcal{N}}(\widehat{\bf x}_{0},\widehat{\bf P}_{0})$ is assumed. By considering the forward sub-model with a frame of light fields ${\bf h}_{k}$ constructed as $p({\bf z}_{k}\mid{\bf x})={\mathcal{N}}(\beta{\bf h}_{k}^{\top}{\bf x},R_{k})$, the posterior PDF after the $k$-th measurement is obtained as
@@
Here a Gaussian approximation $n_{k}\sim{\mathcal{N}}(0,R_{k})$ for the noise model is adopted, where $R_{k}=\beta{\bf h}_{k}^{\top}\widehat{\bf x}_{k-1}$ results from the photon noise assumption, with the current estimate $\widehat{\bf x}_{k-1}$ used to approximate the unknown true $\bf x$}. Under these settings, the posterior statistics ${\widehat{\bf x}}_{k}$, $\widehat{\bf P}_{k}$ are obtained via recursive expressions
@@ 2.1.2 Theoretical objective of AIME
where ${\mathcal{C}}({\bf h}_{k})$ denotes the practical constraints on the encoding ${\bf H}_{k}$. We would like to note that, although derived from an estimation-theoretic perspective, the CRB-based criterion quantifies the same objective of posterior uncertainty reduction as the mutual-information-based formulation. For convenience, this formulation is hereafter referred to as AIME-CRB.
@@ 2.2 Information-optimal encoding in an ideal energy-constrained regime
Figure 2: Illustration of the theoretical result of AIME patterns, each pattern is normalized to its maximum value for visualization. Figures a1, a2 and b1, b2 show the AIME results of two different imaging objects. a1 and b1 are patterns obtained by analytical solution of Eq. (12), each of which show exactly a scanning point; and the position of the point is marked with a purple arrow and box. a2 and b2 are patterns obtained by minimizing the mean CRB, which gradually become a small-size spot as the number of measurements increases; and the spot region in each pattern is displayed in an enlarged view. The first Hermite-Gaussian pattern is employed as the first frame of encoding pattern here.
Based on the information-theoretic objective derived in the previous subsection, we first examine AIME in an idealized yet physically meaningful encoding regime. Specifically, we consider a total-energy constraint on the encoding light fields, which captures a fundamental limitation in optical imaging systems. This regime serves as a benchmark scenario for revealing the intrinsic structure of information-optimal encoding. Under this constraint, the optimal encoding admits a closed-form analytical solution, as established in the following theorem.
@@
This Theorem establishes that, in the ideal energy-constrained regime, information-optimal encoding necessarily reduces to an adaptive point-wise scanning strategy. Importantly, this point-wise structure does not imply a predetermined or fixed scanning order, but instead, varies according to the evolving posterior statistics. For the CRB-based objective shown as Eq. (9), although a closed-form analytical solution of the optimal encoding is difficult to derive, numerical simulations indicate a consistent structural evolution behavior: the optimal encoding initially appears as a diffuse illumination pattern and gradually concentrates into a localized spot as the number of acquired measurements increases.
@@
Figure 3: Experimental comparison of imaging results of AIME and traditional imaging. a: Representative reconstructed images obtained by traditional imaging and AIME under sampling ratios (SRs) of 50% and 100%. Results are shown for two signal-to-noise ratio (SNR) levels, as indicated on the left. The reference image is shown on the right, which is obtained by full-sampling point-wise scanning under a significantly high SNR. b: Quantitative evaluation of imaging quality in terms of PSNR and SSIM as functions of the sampling ratio for the case of ${\rm SNR}=8.74~{\rm dB}$. The horizontal gray lines indicate the performance of traditional full-sampling (100% SR) imaging for reference. c: Corresponding PSNR and SSIM curves for the case of ${\rm SNR}=6.86~{\rm dB}$.
@@
Figure 2 provides an illustrative visualization of the AIME encoding patterns, where panels a and b correspond to two representative imaging objects. To ensure a fair comparison, the first encoding pattern is identically initialized as the first Hermite–Gaussian mode [huang2024compressed] for all cases. The results visually corroborate the theoretical prediction that information-optimal encoding evolves toward an adaptive point-wise structure under the total-energy constraint, which fundamentally differs from conventional raster-scan imaging in that the scanning order is not fixed but adaptively determined by the statistical structure of the object. A quantitative experimental comparison with conventional full-sampling point-wise scanning under identical total-energy and noise conditions is shown in Fig. 3. Here, AIME consistently achieves higher reconstruction quality than traditional imaging, particularly in low-SNR regimes. In specific, with approximately $50\%$ sampling, AIME yields a visually recognizable reconstruction at a detection SNR of $8.47$ dB, whereas conventional imaging with $100\%$ sampling fails to recover the object reliably. Even when both methods operate at full sampling, AIME produces noticeably improved image quality, with an increase of about $0.2$ in SSIM. As the detection SNR further decreases, the object becomes nearly indistinguishable in conventional imaging, while AIME remains capable of providing a coarse yet informative reconstruction. Details of the experimental setting for this comparison is presented in Materials and Methods. These results demonstrate that the adaptive encoding strategy enables more efficient information utilization than fixed raster-scan acquisition in the ideal energy-constrained regime.
@@
An alternative intuition for the emergence of point-wise optimal encoding can be drawn from CS. It is well established in CS theory that sparse signals can be efficiently acquired through incoherent measurements in a conjugate domain, for example, spatially sparse signals can be recovered from partial Fourier measurements [candes2006robust]. In the present imaging scenario, natural scenes are often sparse in transform domains such as the DCT domain. From this perspective, concentrating measurements on a limited set of spatial locations can be viewed as an intuitive counterpart to selectively probing sparse degrees of freedom in a conjugate domain, offering a qualitative analogy to the point-wise scanning behavior predicted by the information-theoretic analysis.
@@ 2.3 Information-optimal encoding under realistic modulation constraints
Figure 4: Simulation results for illustrating the evolution process of the proposed strategy on two different imaging scenes. (a1) and (b1) show the evolution of estimated images and corresponding optimized encoding patterns with sampling ratio increasing for two imaging scenes, respectively; (a2) and (b2) show evolution of PSNR indicators with the increase in sampling; (a3) and (b3) show evolution of estimated PDFs of two selected pixels marked on each, respectively.
@@
which reflects realistic hardware limitations such as the finite modulation depth of spatial light modulators or digital micromirror devices.
@@
Under such constraints, the closed-form analytical solution derived above no longer applies, and the information-optimal encoding is therefore obtained numerically by solving the resulting constrained optimization problem. Figure 4 provides an illustrative visualization of the joint evolution of the reconstructed image and the optimized encoding patterns under the proposed strategy. Two different imaging scenes, namely cameraman and pepper, are considered, and the peak signal-to-noise ratio (PSNR) is used to quantify the imaging quality. Figures 4(a1) and (b1) show the reconstructed images together with the corresponding optimized encoding patterns at different sampling ratios (SRs), where the selected SRs correspond to the marked points in Figs. 4(a2) and (b2), respectively. The optimized encoding patterns exhibit a clear coarse-to-fine evolution as the SR increases. At low SRs, the patterns consist of large-scale speckle-like structures, whereas at higher SRs, finer spatial features gradually emerge. This behavior is consistent with the gradual refinement of the posterior distribution in the Bayesian filtering process, where early measurements focus on coarse information content and subsequent measurements concentrate on increasingly localized spatial details. The corresponding PSNR curves shown in Figs. 4 (a2) and (b2) demonstrate a rapid improvement in reconstruction quality at the early stage of sampling, indicating the high information efficiency of the proposed strategy under limited measurements. To further illustrate the accuracy and uncertainty evolution of the reconstruction, the posterior probability density functions (PDFs) of two representative pixel locations (marked by red and blue crosses in the reconstructed images) are plotted in Figs. 4 (a3) and (b3). As SR increases, the posterior PDFs become progressively sharper, reflecting the systematic reduction of estimation uncertainty and the convergence of pixel-wise intensity estimates. By comparing the results obtained for different imaging scenes, it is evident that the optimized encoding patterns are also strongly scene-dependent under the bounded-amplitude constraint. This observation is fully consistent with the theoretical formulation of AIME and highlights its ability to adaptively allocate measurements according to the estimated statistical state of the imaging object.
@@
Figure 5: Experimental imaging results for comparison between AIME and some typical encoding patterns, a1, a2, and a3 show three different representative cases. For AIME, encoding strategies under both total-energy and amplitude constraints are evaluated, and shown with suffix $\ell_{1}$’ and $ell_{i}nfty$’, respectively. The SR and SNR for these simulations are: a1 SNR around 18 dB, and 25 % SR; a2 SNR around 18.5 dB, and 37.5 % SR; a3 SNR around 20 dB, and 37.5 % SR. Corresponding PSNR and SSIM indicators of imaging results as SR increases are shown in b1, b2, and b3, respectively.
@@
We further compare the proposed AIME strategy with several existing encoding patterns in GI, including pseudo-thermal random patterns, Walsh–Hadamard patterns, and Hermite–Gaussian patterns [huang2024compressed]. For all encoding strategies, the same Bayesian filtering framework is employed for image estimation to ensure a fair comparison. Figure 5 summarizes the imaging results for three representative scenes acquired under different sampling ratios (SRs) and signal-to-noise ratios (SNRs) (Fig. 5 a1–a3). For all tested scenes, both AIME-based strategies consistently yield superior reconstruction quality compared with other encoding patterns, as evidenced by clearer structural details and significantly reduced noise artifacts. The corresponding PSNR and SSIM values are reported under each reconstructed image, where AIME achieves the highest quantitative performance across all cases. The evolution of PSNR and SSIM as functions of the sampling ratio is further plotted in Figs. 5 b1–b3 and c1–c3, respectively. These results reveal that AIME exhibits a pronounced advantage at early measurement stages, where substantial image quality improvements are already achieved at low sampling ratios. As the number of measurements increases, AIME continues to provide steady performance gains, whereas the imaging quality obtained with other encoding patterns rapidly saturates. This behavior demonstrates the strong compressed-sampling capability of AIME and its effectiveness in allocating measurements to the most informative degrees of freedom.
@@
To further quantify the information acquisition capability of different encoding strategies, we explicitly compute the cumulative mutual information and Fisher information accumulated in the detection signals. Specifically, the cumulative mutual information is calculated as $\sum_{k}\frac{1}{2}\ln\left(1+\frac{{{\bf H}_{k}\widehat{\bf P}_{k-1}{\bf H}_{k}^{\top}}}{{R_{k}}}\right)$, while the cumulative Fisher information is evaluated by ${\rm Tr}(\widehat{\bf P}_{k}^{-1})$. The corresponding results is shown in Fig. 5 d1–d3, closely follow the trends observed in the image quality metrics. Across different SRs and noise levels, AIME consistently enables the detection signals to acquire substantially more information about the imaging object than other comparison encoding schemes. This strong agreement between information-theoretic measures and reconstruction quality confirms that the superior imaging performance of AIME is a direct consequence of its enhanced information acquisition capability, rather than heuristic effects.

### v2

@@ Abstract
Ghost imaging (GI) has demonstrated diverse imaging capabilities arising from its encoding-decoding-based computational imaging mechanism. Accordingly, information-theoretic studies have emerged as a promising direction for characterizing the fundamental performance limits of GI and related computational imaging paradigms. Within the information-theoretic framework, the fundamental question of information-optimal encoding design in GI remains largely unexplored, primarily due to the intractability of the prior probability density function (PDF) of an unknown scene. Here, by leveraging the ability of recursively estimating the PDF of the object to be imaged via Bayesian filtering, we establish an adaptive information-maximization encoding (AIME) design framework. Based on the adaptively estimated posterior PDF from previously acquired measurements, the expected information gain of subsequent detections is evaluated and maximized to design the corresponding encoding patterns in a closed-loop manner. Under this framework, the theoretical form of the information-optimal encoding under the energy-constrained regime is analytically derived. Corresponding experimental results show that, GI systems employing information-optimal encoding achieve markedly improved imaging performance compared with conventional fixed point-to-point imaging without relying on additional heuristic regularization, particularly in low signal-to-noise ratio regimes. Moreover, the proposed AIME strategy is compared with existing encoding strategies under a representative experimental constraint, which consistently shows that it enables significantly enhanced information acquisition capability compared with existing encoding strategies, leading to substantially improved imaging quality. These results establish a principled information-theoretic foundation for optimal encoding design in computational imaging paradigms, provided that the forward model can be accurately characterized.
@@ 1 Introduction
Ghost imaging (GI) [moreau2018ghost] has demonstrated a broad range of capabilities, including acquisition of high-dimensional (H-D) information through low-dimensional detection [eshun20253d], enhancement of two-dimensional resolution via H-D light-field information [tong2020breaking, tong2025single], non-locally decoupling detection efficiency and imaging resolution [yuan2021single], and reduction of the unnecessary detection redundancy [hu2019optimization]. These advantages have enabled its diverse applications in imaging LiDAR [deng2017performance, pan2021micro], single-shot H-D imaging cameras [liu2020spectral, chu2021spectral], super-resolution microscopy [li2019single, chen2025multicolor], low-dose and multi-dimensional radiographic imaging [zhang2018tabletop, he2021single], imaging in special band [stantchev2020real, bogdanov2025ghost], and task-oriented imaging [wu2025image, abbas2025target]. From the perspective of imaging information acquisition [barrett2013foundations, shensheng2022review], GI can be interpreted as an encoding-decoding-based computational imaging paradigm [hu2022ghost], where the imaging object is encoded by randomly fluctuating light fields, and the desired image is decoded through high-order light-field correlation. This formulation naturally places GI within an information-theoretic framework, under which analyses of information acquisition efficiency and fundamental performance are in principle attainable [li2017negative, hu2021correspondence, zhou2024theoretical].
@@
In this Article, we explicitly investigate the possibility to adaptively establish the information-optimal encoding for GI, and propose a general adaptive information-maximization encoding (AIME) framework. It enables quantitative evaluation of the information content acquired in the encoding-detection process, and facilitates the design of optimal encoding under practical constraints through information maximization. Theoretical analysis and experimental results consistently reveal that the form of information-optimal encoding is intrinsically governed by the underlying physical constraints. Under an ideal total-energy constraint corresponding to fixed illumination energy, the information-maximization encoding converges to a point-like illumination pattern, leading to an adaptive scanning strategy whose order is determined by the evolving object statistics rather than by a predetermined sequence. Owing to this adaptive design, AIME achieves superior imaging performance compared to conventional “fixed point-to-point” imaging modes without relying on additional heuristic regularization, particularly in low SNR regimes. Under realistic modulation constraints imposed by experimental hardware, although the optimal encoding no longer reduces to a point-like structure, it still emerges as an object-dependent adaptive pattern. In this regime, comparisons with representative encoding strategies show that AIME consistently provides improved imaging quality across a wide range of sampling ratio (SR) and SNR conditions. Further evaluation of the acquired information of different encoding strategies shows strong agreement with image quality, indicating that the performance gain originates from more efficient information acquisition rather than heuristic estimation effects. The proposed framework establishes a principled information-theoretic foundation for characterizing and approaching the fundamental performance of computational imaging paradigms.
@@ 2.1 Principle of AIME for GI
We formulate AIME as a closed-loop information-theoretic framework for GI, in which the encoding patterns are treated as variables to be optimized according to the evolving statistical state of the estimated imaging object. As illustrated in Fig. 1 b, at each measurement step, the statistical state is represented by a posterior PDF inferred from previously detected signals, based on which the expected information gain of a candidate encoding is evaluated. The subsequent encoding is then determined by maximizing the information criterion under explicit physical followed by a recursive update of the posterior PDF after measurement. Within this formulation, encoding, detection, and statistical inference in GI are integrated into a unified framework, enabling information-optimal encoding to be defined and implemented sequentially without requiring special assumption on prior knowledge of the object statistics.
@@ 2.1.1 Bayesian formulation for recursive posterior PDF tracking
To enable information-theoretic evaluation and optimization of encoding, the statistical state of the imaging object is characterized by a posterior PDF that is updated recursively as measurements are acquired. To this end, the forward encoding-detection process of GI is formulated to construct the measurement-likelihood function $p({\bf z}\mid{\bf x})$, specifically as [hu2022ghost, liu2025comprehensive]
@@
Then, an initial Gaussian distribution $p({\bf x}\mid{\bf Z}_{0})=p_{0}({\bf x})={\mathcal{N}}(\widehat{\bf x}_{0},\widehat{\bf P}_{0})$ is assumed by incorporating statistics of imaging scenes in the frequency domain [torralba2003statistics](see Methods and Materials for more details). By considering the forward sub-model with a frame of light fields ${\bf h}_{k}$ constructed as $p({\bf z}_{k}\mid{\bf x})={\mathcal{N}}(\beta{\bf h}_{k}^{\top}{\bf x},R_{k})$, the posterior PDF after the $k$-th measurement is obtained as
@@
Here a Gaussian approximation $n_{k}\sim{\mathcal{N}}(0,R_{k})$ for the noise model is adopted, where $R_{k}=\omega^{2}{\bf h}_{k}^{\top}\widehat{\bf x}_{k-1}$ results from the photon noise assumption [liu2025comprehensive], with the current estimate $\widehat{\bf x}_{k-1}$ used to approximate the unknown true $\bf x$}. Under these settings, the posterior statistics ${\widehat{\bf x}}_{k}$, $\widehat{\bf P}_{k}$ are obtained via recursive expressions
@@
Figure 1: a. Schematic illustration of the ghost imaging (GI) system based on a spatial light modulator. b. Overview of the proposed adaptive information-maximization encoding (AIME) framework for GI. The optimal encoding is designed by maximizing the information contained in the detection signal of the object to be imaged. The information is calculated based on the forward detection model and posterior PDFs.
@@ 2.1.2 Theoretical objective of AIME
    For convenience, this formulation is hereafter referred to as AIME-CRB.
@@
where ${\mathcal{C}}({\bf h}_{k})$ denotes the practical constraints on the encoding ${\bf H}_{k}$. We would like to note that, although derived from an estimation-theoretic perspective, the CRB-based criterion quantifies the same objective of posterior uncertainty reduction as the mutual-information-based formulation.
@@ 2.2 Information-optimal encoding in an ideal energy-constrained regime
Figure 2: Illustration of the theoretical result of AIME-$\ell_{1}$ patterns, each pattern is normalized to its maximum value for visualization. Figures a1, a2 and b1, b2 respectively show AIME results of two different imaging objects. a1 and b1 are patterns obtained by analytical solution of Eq. (12), each of which show exactly a scanning point; and the position of the point is marked with a purple arrow and box. a2 and b2 are patterns obtained by minimizing the mean CRB, which gradually become a small-size spot as the number of measurements increases; and the spot region in each pattern is displayed in an enlarged view. The first Hermite Gaussian pattern is employed as the first frame of encoding pattern here for all cases.
Based on the information-theoretic objective derived in the previous subsection, we first examine AIME in an idealized yet physically meaningful encoding regime. Specifically, we consider a total-energy constraint on the encoding light fields, which captures a fundamental limitation in optical imaging systems. Since the total energy of encoding ${\bf h}_{k}$ can be represented by its $\ell_{1}$ norm, i.e., $\sum_{j}{{\bf h}_{k}}_{j}=\|{\bf h}_{k}\|_{1}$, Aime under this constraint is denoted with suffix $\ell_{1}$”. In this regime, the optimal encoding admits a closed-form analytical solution, as established in the following theorem.
@@
This Theorem establishes that, in the ideal energy-constrained regime, information-optimal encoding necessarily reduces to an adaptive point-wise scanning strategy. Importantly, this point-wise structure does not imply a predetermined or fixed scanning order, but instead, varies according to the evolving posterior statistics. For the CRB-based objective shown as Eq. (9), although a closed-form analytical solution of the optimal encoding is difficult to derive, simulation and experimental results indicate a consistent structural evolution behavior: the optimal encoding initially appears as a diffuse illumination pattern and gradually concentrates into a localized spot as the number of acquired measurements increases.
@@
Figure 3: Experimental comparison of imaging results of GI with AIME-$\ell_{1}$ and conventional imaging. a: Representative imaging results obtained by GI AIME-$\ell_{1}$ and conventional imaging under sampling ratios (SRs) of 50% and 100%. Results are shown for two signal-to-noise ratio (SNR) levels, as indicated on the left. The reference image is shown on the right, which is obtained by full-sampling point-wise scanning under a significantly high SNR. b: Quantitative evaluation of imaging quality in terms of PSNR and SSIM as functions of the sampling ratio for the case of ${\rm SNR}=7.03~{\rm dB}$. The horizontal gray lines indicate the performance of conventional full-sampling (100% SR) imaging for comparison. c: Corresponding PSNR and SSIM curves for the case of ${\rm SNR}=4.85~{\rm dB}$. The horizontal gray lines again indicate the performance of conventional full-sampling (100% SR) imaging for comparison.
@@
Figure 2 provides an illustrative visualization of the AIME-$\ell_{1}$ encoding patterns, where panels a and b correspond to two representative imaging objects. To ensure a fair comparison, the first encoding pattern is identically initialized as the first Hermite Gaussian mode [huang2024compressed] for all cases. The results visually corroborate the theoretical prediction that information-optimal encoding evolves toward an adaptive point-wise structure under the total-energy constraint, which fundamentally differs from conventional raster-scan imaging in that the scanning order is not fixed but adaptively determined by the statistical structure of the object. This difference between AIME-$\ell_{1}$ and conventional imaging lead to an improved imaging performance. To show that, A quantitative experimental comparison under identical total energy and noise conditions is performed. As shown in Fig. 3, the experimental results indicate that AIME consistently achieves higher imaging quality than traditional imaging, particularly in low-SNR regimes. In specific, with approximately 50% sampling, AIME yields a visually recognizable imaging result at a detection SNR of 7.03 dB, whereas conventional imaging with 100% sampling fails to recover the object reliably. When both methods operate at full sampling, AIME produces noticeably improved image quality, with an increase of about 0.2 in SSIM. As the detection SNR further decreases to 4.85 dB, the object becomes nearly indistinguishable in conventional imaging, while AIME remains capable of providing a coarse yet informative image. Details of the experimental setting for this comparison is presented in Materials and Methods, and more experimental comparison results on representative scenes are provided in the supplementary materials. These results demonstrate that the AIME strategy enables more efficient information utilization than fixed raster-scan acquisition in the ideal energy-constrained regime.
@@
An alternative intuition for the emergence of point-wise optimal encoding can be drawn from CS. It is well established in CS theory that sparse signals can be efficiently acquired through incoherent measurements in a conjugate domain, for example, spatially sparse signals can be recovered from partial Fourier measurements [candes2006robust]. In the present imaging scenario, natural scenes are often sparse in transform domains such as the discrete cosine transform (DCT) domain. From this perspective, concentrating measurements on a limited set of spatial locations can be viewed as an intuitive counterpart to selectively probing sparse degrees of freedom in a conjugate domain, offering a qualitative analogy to the point-wise scanning behavior predicted by the information-theoretic analysis.
@@ 2.3 Information-optimal encoding under realistic modulation constraints
which reflects realistic hardware limitations such as the finite modulation depth of spatial light modulators or digital micromirror devices. To differentiate with $\ell_{1}, AIME$ under this constraint is denoted with suffix $\ell_{\infty}$”, since the maximal element of non-negative encoding ${\bf h}_{k}$ can be represented by its $\ell_{\infty}$ norm.
@@
Under such constraints, the closed-form analytical solution derived above no longer applies, and the information-optimal encoding is therefore obtained numerically by solving the resulting constrained optimization problem. Figure 4 provides an illustrative visualization of the joint evolution of imaging results and the optimized encoding patterns under the proposed strategy. Two different imaging scenes, namely cameraman and pepper, are considered, and the peak signal-to-noise ratio (PSNR) is used to quantify the imaging quality. Figures 4 (a1) and (b1) show imaging results together with the corresponding optimized encoding patterns at different sampling ratios (SRs), where the selected SRs correspond to the marked points in Figs. 4 (a2) and (b2), respectively. The optimized encoding patterns exhibit a clear coarse-to-fine evolution as the SR increases. At low SRs, the patterns consist of large-scale speckle-like structures, whereas at higher SRs, finer spatial features gradually emerge. This behavior is consistent with the gradual refinement of the posterior distribution in the Bayesian filtering process, where early measurements focus on coarse information content and subsequent measurements concentrate on increasingly localized spatial details. The corresponding PSNR curves shown in Figs. 4 (a2) and (b2) demonstrate a rapid improvement in reconstruction quality at the early stage of sampling, indicating the high information efficiency of the proposed strategy under limited measurements. To further illustrate the accuracy and uncertainty evolution of the reconstruction, the posterior probability density functions (PDFs) of two representative pixel locations (marked by red and blue crosses in the estimated images) are plotted in Figs. 4 (a3) and (b3). As SR increases, the posterior PDFs become progressively sharper, reflecting the systematic reduction of estimation uncertainty and the convergence of pixel-wise intensity estimates. By comparing the results obtained for different imaging scenes, it is evident that the optimized encoding patterns are also strongly scene-dependent under the bounded-amplitude constraint. This observation is fully consistent with the theoretical formulation of AIME and highlights its ability to adaptively allocate measurements according to the estimated statistical state of the imaging object.
Figure 4: Simulation results for illustrating the evolution process of the proposed strategy on two different imaging scenes. a1 and b1 show the evolution of estimated images and corresponding optimized encoding patterns with sampling ratio increasing for two imaging scenes, respectively; a2 and b2 show evolution of PSNR indicators with the increase in sampling; a3 and b3 show evolution of estimated PDFs of two selected pixels as marked on the imaging results, respectively.
@@
Figure 5: Experimental imaging results for comparison between AIME-$\ell_{\infty}$ and some typical encoding patterns. Results of AIME-$\ell_{1}$ that obtained under the total-energy constraint are also shown here to serve as a fundamental-limit reference. a1, a2, and a3 show three different representative cases. The SR and SNR for them are: a1 SNR around 18 dB, and 25 % SR; a2 SNR around 18.5 dB, and 37.5 % SR; a3 SNR around 20 dB, and 37.5 % SR. PSNR and SSIM indicators of imaging results corresponding to three scenes as SR increases are shown in b1, c1; b2, c2, and b3, c3, respectively. To further quantify and confirm the information acquisition enhancement, the acquired information of different encoding strategies is calculated, and shown in d1, d2, and d3.
@@
To evaluate the practical effectiveness of AIME-$\ell_{\infty}$ under such realistic experimental constraints, we compare it with several representative existing encoding patterns in GI, including pseudo-thermal random patterns, Walsh-Hadamard patterns, and Hermite Gaussian patterns [huang2024compressed]. For theoretical reference, we also include the ideal total-energy-constrained AIME-$\ell_{1}$ results, which serves as a theoretical performance limit without experimental modulation constraints. All results are obtained by employing the same Bayesian filtering framework to ensure a fair comparison that isolates the effect of encoding design. Figure 5 summarizes the results of three representative scenes acquired under different sampling ratios (SRs) and signal-to-noise ratios (SNRs). The imaging results are shown as Figs. 5 a1-a3, respectively. For all tested scenes, both AIME-based strategies (i.e., AIME-MI and AIMe-CRB) consistently yield superior imaging quality compared with other encoding patterns under the same modulation constraints, as evidenced by clearer structural details and significantly reduced noise artifacts. The corresponding PSNR and SSIM values are reported under each image, which confirm that AIME achieves much better quantitative performance across all conditions. At the same time, a performance gap between AIME-$\ell_{\infty}$ and AIME-$\ell_{1}$ theoretical reference can be observed, suggesting the loss of achievable information imposed by hardware constraints on the encoding space. The evolution of PSNR and SSIM as functions of the SR is further plotted in Figs. 5 b1-b3 and c1-c3, respectively. These results reveal that AIME-$\ell_{\infty}$ exhibits a pronounced advantage at early measurement stages, where substantial imaging quality improvements are already achieved at low SRs. As the number of measurements increases, AIME-$\ell_{\infty}$ continues to provide steady performance gains, whereas the imaging quality obtained with other encoding patterns rapidly saturates. This behavior demonstrates the strong compressed-sampling capability of AIME and its effectiveness in allocating the most informative measurements. The gap between AIME-$\ell_{\infty}$ and AIME-$\ell_{1}$ can be also reflected from imaging quality curves, which quantifies the performance loss induced by realistic modulation constraints.
@@
To further quantify the information acquisition capability of different encoding strategies, we explicitly compute the cumulative mutual information accumulated from obtained detection signals. Specifically, the cumulative mutual information is calculated as $\sum_{k}\frac{1}{2}\ln\left(1+\frac{{{\bf H}_{k}\widehat{\bf P}_{k-1}{\bf H}_{k}^{\top}}}{{R_{k}}}\right)$ and plotted on the logarithmic scale. The corresponding results shown in Figs. 5 d1-d3 represent evolution of the acquired amount of information about the unknown imaging scene during the detection process. It can be found that, across different SRs and noise levels, AIME consistently enables the detection signals to acquire substantially more information about the imaging object than other comparison encoding strategies. Also, the separation between AIME-$\ell_{\infty}$ and AIME-$\ell_{1}$ curves quantifies the information loss due to realistic modulation constraints, and mirrors the gap observed in reconstruction quality. Moreover, the ordering of the information curves agrees with the ordering of reconstruction quality, establishing a direct correspondence between information acquisition and imaging performance. This agreement confirms that the superiority of AIME arises from principled information maximization rather than other heuristic effects.